Hey, could someone help me work out this physics-electricity problem? And would you write out your steps please? that would be a big help. thanks!


Two charges of +4.0 nC and -1.0 nC are fixed to a baseline at a separation of 1.0 m. Where on the baseline should a third charge of +2.0 nC be placed if it is to experience

zero net electric force?
Answer #1

Let d is distance from 2 to 4 where 2nC experiences 0 Force.


F(2,4) = F(-1,2)

F = kq1q2/r^2

F(2,4) = k*8*10^-19 / d^2

F(-1,2) = k*2*10^-19 / (1-d)^2

8/d^2 = 2/(1-d)^2

d = 2/3 m = 0.66m


so, it should be placed 0.66 m from 4nC